OpenStudy (anonymous):

how do you graph M(x) = 2|x-3|-4 ?

OpenStudy (nikita2):

http://www.wolframalpha.com/input/?i=2 |x-3|-4

OpenStudy (anonymous):

M(x)=2x-6-4 M(x)=2x-10 let m(x)=y=0 -2x=-10 x=5 (5;0) then let x=0 m(x)=-10 rem.m(x)=y (0;-10)